How much work does the electric field do in moving a proton from a point with a potential of +140 V to a point where it is -60 V ?

A.Express your answer in joules.

b. Express your answer in electron volts.
Answer #1

According to the question we know that

∆U = q∆V = (1.6 × 10−19 C) * ( -60 V - 140 V ) = -3.2 * 10 ^ -17 Joule

so W = 3.2 * 10 ^ -17 Joule ==============PART A)

PART B)

W = 3.2 * 10 ^ -17 Joule / 1.6 × 10−19 J = 200 eV======PART B)